# Alessandro Gambetti1, Qiwei Han1, Maximilian Kaiser2 0000-0003-3389-3784 0000-0002-6044-4530 0009-0007-4329-161X 1Nova School of Business and Economics, Carcavelos, Portugal

### v1

@@ Abstract
Index Terms:
Product Categorization, Multi-Modal Learning, Hierarchical Classification, E-Commerce, Product Recategorization
@@ I Introduction
In the ever-evolving e-commerce landscape, effective product organization is essential for accurate listings, discoverability, operational efficiency, and marketing strategies such as SEO [1]. From a customer perspective, clear taxonomies enable intuitive navigation and improve satisfaction and loyalty [2]. Large platforms rely on automated categorization pipelines, while smaller retailers often depend on manual processes that are time-consuming, error-prone, and costly to scale [3, 4]. As product volumes and diversity continue to expand, scalable automated categorization has become an industrial necessity rather than an optional feature.
Despite its importance, automated categorization remains challenging. Smaller retailers often lack the resources and refined taxonomies of large platforms such as Amazon. Skewed product distributions hinder robust training for long-tail categories [5], and inconsistent category standards across platforms [6, 7] complicate cross-platform analytics. Metadata itself is noisy, such as titles and descriptions vary widely in length, detail, and even language [8], making textual signals unreliable without complementary modalities. Proprietary taxonomies further exacerbate these issues, differing in depth and granularity (e.g., Google [9], Amazon [10]). As shown in Figure 1, the Google Product Taxonomy limits the Shoes category to two levels, while other apparel categories extend to five. Such inconsistencies reduce the effectiveness of downstream applications: recommendation engines cannot distinguish between Sneakers and Formal Shoes if both are grouped under a shallow node. From an industrial perspective, this lack of granularity limits personalization, inventory analytics, and demand forecasting. Existing studies rarely address these platform-specific inconsistencies or the dynamic emergence of new categories [11], underscoring the need for adaptive, cross-platform solutions [12].
@@
  • Hierarchical Classification with Industrial Robustness: We employ a hierarchical classification approach incorporating dynamic masking and shared hidden layers [13]. Rather than optimizing only for benchmark metrics, the design ensures taxonomic consistency and avoids invalid predictions—critical for downstream industrial use cases such as inventory management and recommendation systems.
  • Multimodal Architecture with Cost-Aware Fusion: We integrate textual product attributes (titles and brands) with product images using RoBERTa, ViT, and CLIP. Multiple fusion strategies are evaluated to balance performance and inference cost [14], supporting a two-stage inference pipeline: a lightweight RoBERTa model processes the majority of cases, while a multimodal model is selectively applied to low-confidence predictions, reducing infrastructure costs.
@@ II-A Evolution of Product Categorization: From Unimodal to Multimodal Approaches
Early approaches to product categorization were unimodal, relying on either text or images. Text-based methods analyzed titles and descriptions [12, 15], with [16] categorizing products using only short titles. Vision-based methods employed CNNs to detect main products in fashion images [17]. While effective in narrow settings, text-only systems struggled with noisy or incomplete titles, whereas image-only methods were computationally costly for large catalogs. These limitations motivated multimodal approaches such as CDPF++ [18] and subsequent decision-level fusion [1], which showed that augmenting text with visual features outperformed unimodal baselines. Robustness was further improved by incorporating contextual text [19] and addressing labeling inconsistencies with semi-supervised augmentation [6].
Deep learning approaches, from BERT+ViT with cross-modal attention [20] to pragmatic late-fusion strategies [14], further advanced performance. Pre-trained vision–language models like CLIP [21] have been especially impactful: [22] refined CLIP with cross-modality attention, and [23] applied Siamese networks with BERT and ResNet18 for product matching. Large datasets such as Rakuten [24] enabled evaluation at scale, with [3] confirming the industrial viability of late fusion.
Recently, Large Language Models (LLMs) have been explored for taxonomy tasks. [5] proposed a dual-expert framework combining domain-specific fine-tuning with general LLM reasoning. [25] compared GPT-4o and Claude 3.5 Sonnet for zero-shot categorization, while [4] showed that in-context LLMs can surpass supervised models under noisy metadata. [11] introduced an LLM-agnostic multimodal framework designed to enforce taxonomy consistency. However, LLM adoption in production is limited by inference cost, latency, privacy, and weak consistency guarantees. Supervised multimodal systems remain attractive due to their controllable inference footprint and compatibility with industrial infrastructure.
@@ II-B Flat versus Hierarchical Product Categorization
E-commerce catalogs typically employ hierarchical structures (e.g., Google Product Taxonomy [9]), leading to two main strategies: flat classifiers, which assign items directly to leaf nodes [26], and hierarchical classifiers, which traverse taxonomies level by level [27]. Flat classifiers sometimes outperform but ignore parent–child consistency [9], while hierarchical methods enforce taxonomic coherence [28, 29].
Research on improving hierarchical classification spans multiple directions. [10] improved eBay’s taxonomy accuracy with domain-specific features, and [23] used distributional semantics in a two-level ensemble. Deep learning introduced autoencoders for multi-level classification [12], hierarchical BERT-based models [7], and loss functions penalizing taxonomic errors [30]. Practical deployments include CatReComm [29], an interactive category recommendation system. Dynamic masking [13] further strengthened consistency by constraining fine-level predictions to valid children of predicted parent nodes.
@@ III Data
Each product record is associated with metadata variables including URL, GTIN, Title, Brand, Currency, Price, SKU, Offer ID, Batch ID, and Image URL. In addition, all items were enriched with product categories based on the Google Product Taxonomy, a widely used hierarchical classification system in the retail industry. For example, one product might be assigned to the category Apparel & Accessories ¿ Clothing ¿ Traditional & Ceremonial Clothing ¿ Kimonos ¿ Bridal Kimonos, spanning five levels of depth. Such hierarchical labeling provides a structured foundation for classification tasks, but also reveals systemic imbalances and limitations in standard taxonomies (e.g., shallow versus deep category trees).
@@ IV-A Feature Extraction
For textual attributes, both product titles and brands were encoded separately using a 125M parameter pre-trained RoBERTa model [31]. Embeddings were obtained by extracting the last hidden state and computing the mean across tokens, yielding a dense vector representation for each attribute. This design reflects real-world considerations, as brand and title often provide complementary but noisy signals in industrial datasets. Image features were extracted with a ViT-B/16 model [32], chosen for its ability to capture fine-grained fashion attributes such as patterns and cuts. To reduce inference cost, Principal Component Analysis (PCA) was applied to the image embeddings, preserving 90% of the variance while reducing dimensionality. This trade-off between representational fidelity and computational efficiency proved critical for deployment in GPU-constrained environments.
As an alternative, we also employed CLIP (ViT-B/32) [21], a pre-trained vision-language model that aligns text and images in a shared embedding space. CLIP’s parallel encoders were used to separately process titles, brands, and images. Normalized embeddings were extracted for each modality. Unlike RoBERTa+ViT, CLIP performs alignment during pre-training, making it “fusion-free” in principle. However, downstream classification still requires combining these embeddings, so additional fusion mechanisms were evaluated. The use of CLIP highlights a trade-off familiar to practitioners: pretrained multimodal alignment offers strong generalization but can be more costly to fine-tune at industrial scale.
@@ IV-C Hierarchical Model Architecture
A central innovation in this architecture is Dynamic Masking, illustrated in Figure 4. Following [13], dynamic masking constrains predictions at finer levels by filtering out categories inconsistent with the predicted parent. In practice, once a parent category is selected, a binary mask zeros out non-child categories in the Softmax layer of the next level. This ensures all predicted paths are taxonomically valid and reduces classification noise. From an industrial standpoint, dynamic masking enforces structural consistency without requiring costly post-processing, thereby improving both accuracy and interpretability.
@@ IV-E Evaluation Metrics
However, standard flat metrics alone are insufficient to evaluate hierarchical models, as they do not account for partial correctness in taxonomy-structured predictions [27]. For example, predicting “Shoes” instead of “Sneakers” should not be penalized as harshly as predicting “Handbags.” To capture this nuance, we incorporated hierarchy-aware metrics: hierarchical precision (hP), hierarchical recall (hR), and hierarchical F-score (hF) [33]. These metrics explicitly consider taxonomy paths when computing overlaps between predictions and ground truth, aligning better with real-world expectations for classification consistency in industrial settings.
@@ VI Deployment
One of the major deployment challenges stemmed from the inherent heterogeneity of product categories across platforms. Retailers varied substantially in their taxonomy depth and granularity. Direct-to-consumer (DTC) brand sites and specialized fashion retailers often employed highly specific subcategories, while large generalist marketplaces tended to use broader, less detailed categories. Consequently, the scraped metadata used for categorization—primarily derived from site breadcrumbs—displayed a wide range of inconsistent definitions. The same product could appear under divergent paths across domains, creating difficulties for cross-platform alignment. This lack of standardization impeded downstream analytics, such as attributing category-specific sales effects across platforms, and ultimately degraded the accuracy of EURWEB’s sales forecasting models.
@@ VII Discussion and Conclusion
Our experiments confirmed the advantages of multimodality: CLIP-based late fusion achieved top accuracy, while dynamic masking preserved taxonomy consistency by preventing invalid category paths. Just as importantly, results revealed a practical trade-off: complex fusion methods excel with diverse training data, whereas simpler strategies generalize better to unseen platforms. For deployment, this indicates that industrial systems must not only optimize for peak benchmark accuracy but also for robustness under heterogeneous and resource-limited conditions.
The proposed self-supervised recategorization pipeline further addressed structural gaps in existing taxonomies, generating fine-grained subcategories (e.g., within Shoes) with strong purity levels. By reducing dependence on manual relabeling, this method provides a scalable path for platforms to evolve their taxonomies in response to market trends, aligning classification systems with dynamic product landscapes.
The EURWEB deployment demonstrates how research contributions translate into production. A two-stage pipeline—lightweight RoBERTa for the majority of cases and multimodal RoBERTa+ViT for low-confidence instances—balanced inference cost with classification fidelity. This cascading design reduced GPU overhead while improving category standardization, enabling more reliable forecasting and cross-platform analytics. Such architectures show how advanced multimodal models can be distilled into efficient, production-ready solutions.
Beyond immediate deployment, the framework has broader implications for e-commerce platforms of all sizes. Enhanced multimodal hierarchical classification improves product discoverability, search relevance, and personalized recommendations, while granular taxonomies strengthen inventory analytics, marketing optimization, and long-tail product discoverability. Smaller retailers, which often lack resources to build bespoke systems, can benefit from cross-platform models that transfer knowledge distilled from larger platforms. The semi-automated recategorization pipeline further provides a scalable mechanism for adaptive taxonomy maintenance, reducing manual costs and enabling taxonomies to evolve with emerging product trends. Looking forward, promising directions include extending these methods to other verticals such as electronics and groceries, integrating multilingual capabilities for global deployments, and combining SSL-driven recategorization with large language models (LLMs) for taxonomy generation and maintenance. Additional work on extreme class imbalance, refined dynamic masking for multimodal architectures, and clustering methods that incorporate domain constraints will further strengthen robustness and industrial applicability.

### v2

@@ I Introduction
In the ever-evolving e-commerce landscape, effective product organization is essential for accurate listings, discoverability, operational efficiency, and marketing strategies such as SEO [Zahavy2018]. From a customer perspective, clear taxonomies enable intuitive navigation and improve satisfaction and loyalty [girard2002influence]. Large platforms rely on automated categorization pipelines, while smaller retailers often depend on manual processes that are time-consuming, error-prone, and costly to scale [Bi2020, gholamian-etal-2024-llm]. As product volumes and diversity continue to expand, scalable automated categorization has become an industrial necessity rather than an optional feature.
Despite its importance, automated categorization remains challenging. Smaller retailers often lack the resources and refined taxonomies of large platforms such as Amazon. Skewed product distributions hinder robust training for long-tail categories [cheng2024commerce], and inconsistent category standards across platforms [avigdor2023consistent, zhang2021deep] complicate cross-platform analytics. Metadata itself is noisy, such as titles and descriptions vary widely in length, detail, and even language [de2018generating], making textual signals unreliable without complementary modalities. Proprietary taxonomies further exacerbate these issues, differing in depth and granularity (e.g., Google [Krishnan2019], Amazon [Shen2011]). As shown in Figure 1, the Google Product Taxonomy limits the Shoes category to two levels, while other apparel categories extend to five. Such inconsistencies reduce the effectiveness of downstream applications: recommendation engines cannot distinguish between Sneakers and Formal Shoes if both are grouped under a shallow node. From an industrial perspective, this lack of granularity limits personalization, inventory analytics, and demand forecasting. Existing studies rarely address these platform-specific inconsistencies or the dynamic emergence of new categories [chen-etal-2025-leveraging], underscoring the need for adaptive, cross-platform solutions [Cevahir2016].
@@
  • Hierarchical Classification with Industrial Robustness: We employ a hierarchical classification approach incorporating dynamic masking and shared hidden layers [Ozyegen2022]. Rather than optimizing only for benchmark metrics, the design ensures taxonomic consistency and avoids invalid predictions—critical for downstream industrial use cases such as inventory management and recommendation systems.
  • Multimodal Architecture with Cost-Aware Fusion: We integrate textual product attributes (titles and brands) with product images using RoBERTa, ViT, and CLIP. Multiple fusion strategies are evaluated to balance performance and inference cost [Tashu2022], supporting a two-stage inference pipeline: a lightweight RoBERTa model processes the majority of cases, while a multimodal model is selectively applied to low-confidence predictions, reducing infrastructure costs.
@@ II-A Evolution of Product Categorization: From Unimodal to Multimodal Approaches
Early approaches to product categorization were unimodal, relying on either text or images. Text-based methods analyzed titles and descriptions [Cevahir2016, tan2020commerce], with [Kozareva2015] categorizing products using only short titles. Vision-based methods employed CNNs to detect main products in fashion images [Yu2017]. While effective in narrow settings, text-only systems struggled with noisy or incomplete titles, whereas image-only methods were computationally costly for large catalogs. These limitations motivated multimodal approaches such as CDPF++ [Kannan2011] and subsequent decision-level fusion [Zahavy2018], which showed that augmenting text with visual features outperformed unimodal baselines. Robustness was further improved by incorporating contextual text [Kalva2007] and addressing labeling inconsistencies with semi-supervised augmentation [avigdor2023consistent].
Deep learning approaches, from BERT+ViT with cross-modal attention [Chen2021] to pragmatic late-fusion strategies [Tashu2022], further advanced performance. Pre-trained vision–language models like CLIP [radford2021learning] have been especially impactful: [Fu2022] refined CLIP with cross-modality attention, and [Gupta2016] applied Siamese networks with BERT and ResNet18 for product matching. Large datasets such as Rakuten [Amoualian2021] enabled evaluation at scale, with [Bi2020] confirming the industrial viability of late fusion.
Recently, Large Language Models (LLMs) have been explored for taxonomy tasks. [cheng2024commerce] proposed a dual-expert framework combining domain-specific fine-tuning with general LLM reasoning. [roumeliotis2025llms] compared GPT-4o and Claude 3.5 Sonnet for zero-shot categorization, while [gholamian-etal-2024-llm] showed that in-context LLMs can surpass supervised models under noisy metadata. [chen-etal-2025-leveraging] introduced an LLM-agnostic multimodal framework designed to enforce taxonomy consistency. However, LLM adoption in production is limited by inference cost, latency, privacy, and weak consistency guarantees. Supervised multimodal systems remain attractive due to their controllable inference footprint and compatibility with industrial infrastructure.
@@ II-B Flat versus Hierarchical Product Categorization
E-commerce catalogs typically employ hierarchical structures (e.g., Google Product Taxonomy [Krishnan2019]), leading to two main strategies: flat classifiers, which assign items directly to leaf nodes [cotacallapa2024flat], and hierarchical classifiers, which traverse taxonomies level by level [Silla2010]. Flat classifiers sometimes outperform but ignore parent–child consistency [Krishnan2019], while hierarchical methods enforce taxonomic coherence [Bergamaschi2002, Hasson2021].
Research on improving hierarchical classification spans multiple directions. [Shen2011] improved eBay’s taxonomy accuracy with domain-specific features, and [Gupta2016] used distributional semantics in a two-level ensemble. Deep learning introduced autoencoders for multi-level classification [Cevahir2016], hierarchical BERT-based models [zhang2021deep], and loss functions penalizing taxonomic errors [Gao2020]. Practical deployments include CatReComm [Hasson2021], an interactive category recommendation system. Dynamic masking [Ozyegen2022] further strengthened consistency by constraining fine-level predictions to valid children of predicted parent nodes.
@@ III Data
Each product record is associated with metadata variables including URL, GTIN, Title, Brand, Currency, Price, SKU, Offer ID, Batch ID, and Image URL. In addition, all items were enriched with product categories based on the Google Product Taxonomy, a widely used hierarchical classification system in the retail industry. For example, one product might be assigned to the category Apparel & Accessories >> Clothing >> Traditional & Ceremonial Clothing >> Kimonos >> Bridal Kimonos, spanning five levels of depth. Such hierarchical labeling provides a structured foundation for classification tasks, but also reveals systemic imbalances and limitations in standard taxonomies (e.g., shallow versus deep category trees).
@@ IV-A Feature Extraction
For textual attributes, both product titles and brands were encoded separately using a 125M parameter pre-trained RoBERTa model [liu2019roberta]. Embeddings were obtained by extracting the last hidden state and computing the mean across tokens, yielding a dense vector representation for each attribute. This design reflects real-world considerations, as brand and title often provide complementary but noisy signals in industrial datasets. Image features were extracted with a ViT-B/16 model [dosovitskiy2021image], chosen for its ability to capture fine-grained fashion attributes such as patterns and cuts. To reduce inference cost, Principal Component Analysis (PCA) was applied to the image embeddings, preserving 90% of the variance while reducing dimensionality. This trade-off between representational fidelity and computational efficiency proved critical for deployment in GPU-constrained environments.
As an alternative, we also employed CLIP (ViT-B/32) [radford2021learning], a pre-trained vision-language model that aligns text and images in a shared embedding space. CLIP’s parallel encoders were used to separately process titles, brands, and images. Normalized embeddings were extracted for each modality. Unlike RoBERTa+ViT, CLIP performs alignment during pre-training, making it “fusion-free” in principle. However, downstream classification still requires combining these embeddings, so additional fusion mechanisms were evaluated. The use of CLIP highlights a trade-off familiar to practitioners: pretrained multimodal alignment offers strong generalization but can be more costly to fine-tune at industrial scale.
@@ IV-C Hierarchical Model Architecture
A central innovation in this architecture is Dynamic Masking, illustrated in Figure 4. Following [Ozyegen2022], dynamic masking constrains predictions at finer levels by filtering out categories inconsistent with the predicted parent. In practice, once a parent category is selected, a binary mask zeros out non-child categories in the Softmax layer of the next level. This ensures all predicted paths are taxonomically valid and reduces classification noise. From an industrial standpoint, dynamic masking enforces structural consistency without requiring costly post-processing, thereby improving both accuracy and interpretability.
@@ IV-E Evaluation Metrics
However, standard flat metrics alone are insufficient to evaluate hierarchical models, as they do not account for partial correctness in taxonomy-structured predictions [Silla2010]. For example, predicting “Shoes” instead of “Sneakers” should not be penalized as harshly as predicting “Handbags.” To capture this nuance, we incorporated hierarchy-aware metrics: hierarchical precision (hP), hierarchical recall (hR), and hierarchical F-score (hF) [miranda2023hiclass]. These metrics explicitly consider taxonomy paths when computing overlaps between predictions and ground truth, aligning better with real-world expectations for classification consistency in industrial settings.
@@ VI Deployment
One of the major deployment challenges stemmed from the inherent heterogeneity of product categories across platforms. Retailers varied substantially in their taxonomy depth and granularity. Direct-to-consumer (DTC) brand sites and specialized fashion retailers often employed highly specific subcategories, while large generalist marketplaces tended to use broader, less detailed categories. Consequently, the scraped metadata used for categorization primarily derived from site breadcrumbs, displayed a wide range of inconsistent definitions. The same product could appear under divergent paths across domains, creating difficulties for cross-platform alignment. This lack of standardization impeded downstream analytics, such as attributing category-specific sales effects across platforms, and ultimately degraded the accuracy of EURWEB’s sales forecasting models.
@@ VII Discussion and Conclusion
Our experiments confirmed the advantages of multimodality: CLIP-based late fusion achieved top accuracy, while dynamic masking preserved taxonomy consistency by preventing invalid category paths. Just as importantly, results revealed a practical trade-off: complex fusion methods excel with diverse training data, whereas simpler strategies generalize better to unseen platforms. For deployment, this indicates that industrial systems must not only optimize for peak benchmark accuracy but also for robustness under heterogeneous and resource-limited conditions. The proposed self-supervised recategorization pipeline further addressed structural gaps in existing taxonomies, generating fine-grained subcategories (e.g., within Shoes) with strong purity levels. By reducing dependence on manual relabeling, this method provides a scalable path for platforms to evolve their taxonomies in response to market trends, aligning classification systems with dynamic product landscapes.
The EURWEB deployment demonstrates how research contributions translate into production. A two-stage pipeline: lightweight RoBERTa for the majority of cases and multimodal RoBERTa+ViT for low-confidence instances, balanced inference cost with classification fidelity. This cascading design reduced GPU overhead while improving category standardization, enabling more reliable forecasting and cross-platform analytics. Such architectures show how advanced multimodal models can be distilled into efficient, production-ready solutions.
Beyond immediate deployment, the framework has broader implications for e-commerce platforms of all sizes. Enhanced multimodal hierarchical classification improves product discoverability, search relevance, and personalized recommendations, while granular taxonomies strengthen inventory analytics, marketing optimization, and long-tail product discoverability. Smaller retailers, which often lack resources to build bespoke systems, can benefit from cross-platform models that transfer knowledge distilled from larger platforms. The semi-automated recategorization pipeline further provides a scalable mechanism for adaptive taxonomy maintenance, reducing manual costs and enabling taxonomies to evolve with emerging product trends. Looking forward, promising directions include extending these methods to other verticals such as electronics and groceries, integrating multilingual capabilities for global deployments, and combining SSL-driven recategorization with large language models (LLMs) for taxonomy generation and maintenance [cheng2024commerce, chen-etal-2025-leveraging, gholamian-etal-2024-llm]. Additional work on extreme class imbalance, refined dynamic masking for multimodal architectures, and clustering methods that incorporate domain constraints will further strengthen robustness and industrial applicability.